OpenStudy (anonymous):

find the center and radius of the cirlce x^2+y^2-6x+4y+9=0

OpenStudy (anonymous):

complete the square twice \[x^2-6x+y^2+4y=-9\] is a start then for the \(x\) terms you get \[(x-3)^2+y^2+4y=-9+9=0\] can you do it for the \(y\) terms?

OpenStudy (anonymous):

i actually kind of figured out but for the center im hesistating that if its -3 and 2

OpenStudy (anonymous):

you got it backwards, it is \((3,-2)\)

OpenStudy (anonymous):

general form is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

thank you! do you know anything about canonical equation?